### v1

# Axion-like Dark Matter Search with Space-based Gravitational Wave Detectors

Abstract
We propose a novel modification to the optical benches of space-based gravitational wave detectors (SGWDs) to enable the detection of axion-like dark matter (ALDM)-induced birefringence without altering the polarization of inter-spacecraft laser links. Our design introduces an auxiliary interferometer to convert polarization modulation into measurable phase shifts. Analytical expressions for sensitivity to the ALDM-photon coupling are derived for various time-delay interferometry (TDI) combinations. Projected sensitivity curves demonstrate complementary coverage across the ALDM mass range $10^{-19}\sim 10^{-14}\mathrm{eV}$. This approach preserves the original interferometric stability while enabling new physics capabilities for SGWDs.
I Introduction
The endeavor to unravel the nature of dark matter, one of the most profound and enduring mysteries in modern physics Zwicky (1933); Bertone and Tait (2018), has driven scientists to explore a wide range of theoretical candidates, ranging from ultra-light scalar particles to primordial black holes. Among these possibilities, axion-like dark matter (ALDM) has emerged as a particularly compelling candidate Preskill et al. (1983); Abbott and Sikivie (1983); Dine and Fischler (1983); Peccei (2008); Marsh (2016); Adams et al. (2022); O’Hare (2024), distinguished by a unique set of properties that make it an attractive focus for both theoretical and experimental investigations Arvanitaki et al. (2010); Graham et al. (2015); Di Luzio et al. (2020). These particles are postulated to possess an extraordinarily low mass, exhibit exceedingly feeble interactions, and exist in large abundance.
One of the most intriguing aspects of ALDM is its interaction with electromagnetic fields, which gives rise to distinctive signatures that are potentially detectable. A key phenomenon associated with ALDM is the birefringence effect Carroll et al. (1990); Carroll (1998); Harari and Sikivie (1992), which arises from the coupling between the ALDM field and photons. In the presence of an ALDM field, the phase velocities of left-handed and right-handed circularly polarized light differ while the light propagating through space. This difference causes linearly polarized light to undergo a rotation of its polarization plane as it propagates. This effect provides a distinctive and measurable signature of ALDM. By studying this birefringence effect, researchers can probe the strength of the ALDM-photon coupling, offering a powerful tool to test the existence of these elusive particles and constrain their intrinsic properties.
While traditional astrophysical observations, such as cosmic microwave background (CMB) polarization studies Harari and Sikivie (1992); Fedderke et al. (2019); Fujita et al. (2021) and celestial polarimetry di Serego Alighieri et al. (2010); Fujita et al. (2019); Ivanov et al. (2019); Liu et al. (2020); Poddar and Mohanty (2020); Chen et al. (2022); Yuan et al. (2021); Liu et al. (2023), have been employed to search for birefringence effects induced by ALDM, their sensitivity is ultimately limited by systematic uncertainties and astrophysical foregrounds. To achieve greater precision, new experimental approaches leveraging high-precision interferometry have been proposed. In particular, optical cavities and laser interferometers DeRocco and Hook (2018); Obata et al. (2018); Liu et al. (2019); Nagano et al. (2019); Michimura et al. (2020); Heinze et al. (2024), which can measure tiny phase shifts with extraordinary accuracy, present promising avenues for detecting the subtle imprint of ALDM-induced birefringence.
Gravitational wave (GW) detectors are well-developed interferometers for probing the phase signature induced by ALDM. The detection of GW has revolutionized astrophysics since the first observation by the LIGO-Virgo collaboration in 2015 Abbott et al. (2016), opening a novel observational window into the cosmos. Ground-based interferometers, such as LIGO Aasi et al. (2015), Virgo Acernese et al. (2015), and KAGRA Akutsu et al. (2019), now routinely detect high-frequency GWs ($10\,\mathrm{Hz}\sim 10\,\mathrm{kHz}$) originating from compact binary mergers, including black holes and neutron stars, while pulsar timing arrays McLaughlin (2013); Manchester et al. (2013); Kramer and Champion (2013); Manchester (2013) probe the nanohertz regime in search of stochastic signals Agazie et al. (2023); Reardon et al. (2023); Antoniadis et al. (2023); Xu et al. (2023) from supermassive black hole binaries. However, the millihertz frequency band—home to mergers of massive black hole binaries, extreme mass-ratio inspirals, and Galactic compact binaries—remains inaccessible to current ground-based or PTA efforts. This gap will be addressed by future space-based detectors, such as LISA Colpi et al. (2024), Taiji Luo et al. (2020), TianQin Luo et al. (2016), and Big-Bang Observer (BBO) Crowder and Cornish (2005), which utilize laser interferometry between spacecraft separated by millions of kilometers to achieve unprecedented phase sensitivity. These missions employ advanced technologies, including picometer-level metrology, drag-free spacecraft, and time-delay interferometry (TDI) Tinto and Dhurandhar (2021), to suppress noise sources, thereby enabling not only GW astronomy but also synergies with fundamental physics.
Given the exceptional phase sensitivity of SGWDs, typically at the picometer level over million-kilometer baselines, these instruments hold significant potential for detecting the subtle effects induced by the ALDM. Lately, modifications to these SGWDs have been proposed to probe ALDM Gué et al. (2025); Yao et al. (2025). These studies suggest altering the polarization of the laser beam exchanged between the spacecraft (S/C) from linear to circular. With these modificaitons, the phase difference induced by ALDM would be integrated into the phase measurement of SGWDs. However, such modifications may introduce some issues like the susceptibility to roll rotations of the S/C Müller (2017) and the generation of backreflection LIS (2000).
In this study, we propose an alternative approach that preserves linear polarization while adding a dedicated auxiliary interferometer on each optical bench (OB). This supplementary interferometer is designed to convert small polarization rotations into a differential phase signal, that can be measured using the established heterodyne detection schemes already implemented for gravitational wave observations. Our approach offers several advantages, including compatibility with current OB designs, the use of established phase measurement techniques with picometer-level sensitivity, and simultaneous operation with gravitational wave detection.
The paper is organized as follows. In Section II we review the ALDM-induced birefringence effect and its impact on the polarization state. Section III introduces our proposed modification to the optical bench design. Section IV presents a detailed sensitivity analysis of the ALDM-photon coupling. Section V discusses practical considerations. Finally, the conclusion is given in Section VI. Throughout this paper, all equations are expressed in Lorentz–Heaviside units.
II ALDM-induced Birefringence
In this section, we briefly introduce the birefringence effect induced by ALDM and its implications for space-based interferometry. The ALDM-photon interaction is decribed by the Lagrangian Carroll et al. (1990); Carroll and Field (1991); Harari and Sikivie (1992)
$\mathcal{L}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}+\frac{1}{2}\partial^{\mu}a\partial_{\mu}a-\frac{1}{2}m_{a}^{2}a^{2}-\frac{1}{4}g_{a\gamma}aF_{\mu\nu}\tilde{F}^{\mu\nu},$ (1)
where $a$ is the ALDM field with mass $m_{a}$, $g_{a\gamma}$ is the ALDM-photon coupling, and $F^{\mu\nu}$ and $\tilde{F}^{\mu\nu}$ are the electromagnetic stress tensor and its dual, respectively.
While photons propagate through the ALDM field, the dispersion relations for the left and the right circular polarization modes attain opposite corrections due to the temporal and spatial variations of the ALDM field, which can be expressed as
$\omega_{\pm}\simeq k\pm g_{a\gamma}\left(\frac{\partial a}{\partial t}+\nabla a\cdot\bm{\hat{k}}\right)=k\pm g_{a\gamma}n^{\mu}\nabla_{\mu}a,$ (2)
where $n^{\mu}$ denotes the null tangent vector to the photon path. This effect is known as ALDM-induced birefringence.
The differential dispersion relation implies that left and right circularly polarized waves propagate with slightly different phase velocities, leading to a cumulative phase difference between them. For linearly polarized light, which can be decomposed into equal amplitudes of left and right circular polarizations, this phase difference manifests as a rotation of the polarization plane.
The phase difference between the left and right circular polarization can be probed by SGWDs, constituting the central detection principle proposed in Gué et al. (2025); Yao et al. (2025). In their proposals, the linearly polarized light transmitted between S/C is substituted with circularly polarized light, and the phase shift induced by the birefringence effect is integrated into the interferometer signal. However, although the current design of SGWDs utilize linearly polarized laser links, the past designs of SGWDs favored circularly polarized light for the inter-satellite path Müller (2017), and both options were under consideration ESA (2011). It seems that the advantages of linear polarization were deemed more favorable after detailed trade-off studies. Using circular polarization requires converting from linear to circular via quarter-wave plates and then back to linear at the receiving end. Extra conversion steps are sensitive to relative rotations, coating phase shifts, and alignment errors between spacecraft. Circular polarization in phase sensitive ranging yields a susceptibility to roll rotations of the S/C Müller (2017). Besides, optical components, such as quarter-wave plates, are main sources of backreflection, and it is easier to design and to predict the behaviour of the antireflection coatings of the optics working only with linear polarisation LIS (2000). This motivates the exploration of alternative approaches that preserve the primary light path of SGWDs.
For linearly polarized light, the differential shift leads to a rotation of the electric vector position angle, expressed as:
[…]
$=$ $\frac{1}{2}g_{a\gamma}\left[a(x_{\mathrm{obs}}^{\mu})-a(x_{\mathrm{emit}}^{\mu})\right].$
This rotation only depends on the difference in the ALDM field values at emission and observation points, denoted by $x_{\mathrm{emit}}^{\mu}$ and $x_{\mathrm{obs}}^{\mu}$.
Neglecting the back-action term of light on the ALDM, the wave function of the ALDM filed, derived from the equations of motion arising from Eq. (1), can be written as
$a(t,\bm{x})=a_{0}\cos(\omega_{a}t-\bm{k}_{a}\cdot\bm{x}+\Phi),$ (4)
where $a_{0}$ is the field amplitude, related to the dark matter density $\rho_{\mathrm{DM}}=m_{a}^{2}a_{0}^{2}/2\approx 0.4\>\mathrm{GeV\cdot cm^{-3}}$ McMillan (2011), and $\Phi$ denotes an unknown phase. In the non-relativistic limit, the frequency $\omega_{a}$ and wave vector $\bm{k}_{a}$ are given by $\omega_{a}\simeq m_{a}(1+\bm{v}^{2}/2)$ and $\bm{k}_{a}\simeq m_{a}\bm{v}$, respectively, where $\bm{v}$ is the collective velocity of the ALDM with $v\sim 10^{-3}$. When the coherence length of the ALDM field, $\lambda_{a}=2\pi/k_{a}$, significantly exceeds the distance between S/C, denoted as $L$, Eq. (4) can be approximated as
$a(t)\approx\frac{\sqrt{2\rho_{\mathrm{DM}}}}{m_{a}}\cos(m_{a}t+\Phi).$ (5)
This approximation is valid for the majority of the ALDM mass range of interest ($10^{-19}$ to $10^{-14}$ eV), as the corresponding coherence length exceeds the typical arm length of SGWDs, which is on the order of $\sim 10^{9}\>\mathrm{m}$. Under these conditions, the ALDM field can be treated as temporally oscillating but spatially uniform across the detector network.
The expected change in the polarization angle induced by the birefringence effect for a single-link light reads
$\Delta\phi_{a}=$ $\frac{1}{2}g_{a\gamma}\frac{\sqrt{2\rho_{\mathrm{DM}}}}{m_{a}}\left\{\cos(m_{a}t+\Phi)-\cos\left[m_{a}(t-L)+\Phi\right]\right\}$ (6)
$=$ $-g_{a\gamma}\frac{\sqrt{2\rho_{\mathrm{DM}}}}{m_{a}}\sin\frac{m_{a}L}{2}\sin\left(m_{a}t+\Phi-\frac{m_{a}L}{2}\right)$
$\sim$ $-10^{-7}\left(\frac{g_{a\gamma}}{10^{-11}\,\mathrm{GeV}^{-1}}\right)\left(\frac{\rho_{\mathrm{DM}}}{0.4\,\mathrm{GeV}\cdot\mathrm{cm}^{-3}}\right)^{1/2}\left(\frac{10^{-16}\,\mathrm{eV}}{m_{a}}\right)$
×$\times\sin\left(\frac{m_{a}L}{2}\right)\sin\left(m_{a}t+\Phi-\frac{m_{a}L}{2}\right).$
The numerical estimate in the above equation reveals that the birefringence-induced polarization rotation is quite small, approximately $10^{-7}$ radians for typical parameter values. This underscores the challenge of detecting such a signal and motivates the need for highly sensitive interferometric techniques.
Without loss of generality, we suppose that the transmitting laser light from remote S/C is s-polarized with an amplitude $E_{\mathrm{TX}}$. The received light affected by the ALDM oscillation can be written as
$\bm{E}_{\mathrm{RX}}=\cos\Delta\phi_{a}E_{\mathrm{TX}}\hat{\bm{s}}+\sin\Delta\phi_{a}E_{\mathrm{TX}}\hat{\bm{p}},$ (7)
where $\hat{\bm{s}}$ and $\hat{\bm{p}}$ represent the unit vectors in the s-polarization and p-polarization directions, respectively. The presence of the ALDM field induces a small p-polarization component proportional to $\sin\Delta\phi_{a}$, while slightly reducing the s-polarization component proportional to $\cos\Delta\phi_{a}$.
From Eq. (7), the s-polarized component of the received light undergoes periodic modulation at $\mathcal{O}(\Delta\phi_{a}^{2})$ due to the ALDM field. However, this modulation lies below the sensitivity threshold of the current design of SGWDs. A possible approach to make the interferometer sensitive to the birefringence effect is to measure the amplitude of the p-polarized component of the received light, as proposed in studies such as Nagano et al. (2019). Nevertheless, direct detection of the p-polarized signal poses significant challenges in SGWDs, due to the extremely low power of the received light, which is on the order of $\sim 700\,\mathrm{pW}$. Our proposal circumvents this limitation by converting the polarization rotation into a differential phase measurement through the implementation of an auxiliary interferometer.
III Proposed Modification
Figure 1: Schematic of the modified optical bench design. The dotted box highlights the additional ALDM interferometer, phase shifter, and associated optics added to the current OB configuration. The orange beam represents the transmitted light, the green dotted path indicates the received s-polarized light, and the green waving path shows the received p-polarized light generated by ALDM-induced birefringence. PBS: polarizing beam splitter; BS: beam splitter; $\lambda/2$: half-wave plate; PAAM: point-ahead angle mechanism; A,B,C,D: photodector.
In order to measure the ALDM-induced effect without altering the linearly polarized light, a viable solution is to convert power measurement to phase measurement. This can be achieved by introducing an auxiliary interferometer for the p-polarized component of the received light on the OB. In this work, we propose a modification of the OB design to detect the birefringence effect signals.
The main innovation of our approach lies in maintaining the established linear polarization scheme while adding a dedicated detection path for the ALDM-induced p-polarization component. This strategy preserves the advantages of the current SGWD design while enhancing sensitivity to new physics.
The proposed modification to the OB is illustrated in Fig. 1. The optical elements enclosed within the dotted line, including the ALDM interferometer, a phase shifter, and associated mirrors, are added to the OB, while the remaining components represent a simplified version of the current OB design for SGWDs. Polarizing beam splitters (PBS) are employed to separate the transmitted and received light. In the schematic, the transmitted light is depicted in orange, the received s-polarized light is represented by green dotted lines, and the received p-polarized light is illustrated with green wavy lines.
In our proposed design, the PBS directs the p-polarized component, which contains the ALDM signal, into the auxiliary interferometer path, while allowing the s-polarized component to propagate along the standard interferometer path. The auxiliary path includes a phase shifter that enables precise control of the relative phase between the two interferometer outputs, which is crucial for optimizing the detection sensitivity.
For the inter-spacecraft interferometer (ISI), the interference between a local transmitted beam and the received beam results in heterodyne signals at the photodiodes. In the presence of the ALDM field, the received amplitude is modulated by a factor of $\cos\Delta\phi_{a}$, while in the ALDM interferometer, the modulation appears as $\sin\Delta\phi_{a}$. With the auxiliary channel, the birefringence-induced amplitude modulation is converted into a phase shift $\phi_{\mathrm{ALDM}}$, which can be read out by combining the heterodyne outputs of the ISI and the ALDM interferometer appropriately. The phase shift signal is then compared with that from the ISI, $\phi_{\mathrm{ISI}}$, as follows
$\phi_{\mathrm{ALDM}}-\phi_{\mathrm{ISI}}\simeq\Delta\phi_{a},$ (8)
thereby isolating the ALDM contribution. This differential measurement approach offers a significant advantage: common-mode noises, such as laser frequency fluctuations and S/C motion, are largely canceled out, leaving primarily the ALDM-induced signal.
In the ISI, a fraction of the local transmitted beam $\sqrt{2I_{0}}e^{i\omega_{\mathrm{T}}t}$ interferes with the light received from the remote S/C $p_{\mathrm{ISI}}\sqrt{2I_{0}}e^{i(\omega_{R}t+\phi)}$, where $I_{ref}=2I_{0}$ is the power of the local reference, $p_{\mathrm{ISI}}^{2}\sim 10^{-7}$ is the ratio of the received power to the local reference, $\omega_{\mathrm{T}}$ is the angular frequency of the transmitted beam, $\omega_{R}$ is the angular frequency of the received beam, and $\phi$ is the phase of the received wavefront. In the presence of the ALDM field, the power of the received light in the ISI is modulated by the birefringence effect, expressed as
$p_{\mathrm{ISI}}^{\prime}=\cos\Delta\phi_{a}p_{\mathrm{ISI}}.$ (9)
The total field at the photodiodes include both the primary signal and potential noise sources, such as backscattered light, which can introduce spurious signals if not properly mitigated. Its expression can be given by
$E_{A}=\sqrt{2I_{0}}e^{i\omega_{\mathrm{T}}t}\left[1+p_{A}^{\prime}e^{i(\Omega t+\phi+\pi/2)}+u_{A}e^{i(\psi+\pi/2)}\right],$ (10)
$E_{B}=\sqrt{2I_{0}}e^{i\omega_{\mathrm{T}}t}\left[e^{i\pi/2}+p_{B}^{\prime}e^{i(\Omega t+\phi)}+u_{B}e^{i\psi}\right],$ (11)
where the subscripts A and B label the diode ports, $\Omega=\omega_{R}-\omega_{\mathrm{T}}$ is the heterodyne frequency, $p_{\mathrm{ISI}}^{\prime 2}=p_{A}^{\prime 2}+p_{B}^{\prime 2}\ll 1$, the fraction $u_{\mathrm{ISI}}^{2}=u_{A}^{2}+u_{B}^{2}\ll 1$ and phase $\psi$ characterizes the field of the backscattered light. The heterodyne signals at the photodiode A and B are
$S_{A}(t)=2I_{0}\left[1+p_{A}^{\prime 2}+u_{A}^{2}-2u_{A}\sin\psi-2p_{A}^{\prime}\sin(\Omega t+\phi)+2p_{A}^{\prime}u_{A}\cos(\Omega t+\phi-\psi)\right],$ (12)
$S_{B}(t)=2I_{0}\left[1+p_{B}^{\prime 2}+u_{B}^{2}+2u_{B}\sin\psi+2p_{B}^{\prime}\sin(\Omega t+\phi)+2p_{B}^{\prime}u_{B}\cos(\Omega t+\phi-\psi)\right].$ (13)
The photodiode outputs contain the desired heterodyne signal at frequency $\Omega$, along with direct current (DC) terms and beat signals between the received and backscattered light.
The digitized signal difference is
$S_{B}(t)-S_{A}(t)$ $=2I_{0}\left[p_{B}^{\prime 2}-p_{A}^{\prime 2}+u_{B}^{2}-u_{A}^{2}+2(u_{B}+u_{A})\sin\psi\right.$ (14)
$\left.+2(p_{B}^{\prime}+p_{A}^{\prime})\sin(\Omega t+\phi)+2(p_{B}^{\prime}u_{B}-p_{A}^{\prime}u_{A})\cos(\Omega t+\phi-\psi)\right].$
This signal difference is multiplied by a local oscillator with frequency $\Omega$,
[…]
where the integration extends over $n$ oscillator cycles. The integrated output is proportional to the phase difference between the signal and oscillator, where only alternating current (AC) terms contribute. This demodulation process effectively extracts the phase information from the heterodyne beat signal, which constitutes the primary quantity measured in the interferometer.
From the integrated output, the phase of the heterodyne-signal can be recovered as Sasso et al. (2019)
$\phi_{\mathrm{ISI}}\approx\phi+\pi/2+\tilde{u}\cos\psi-\tilde{u}^{2}\sin\psi\cos\psi,$ (16)
with
$\tilde{u}_{\mathrm{ISI}}=\frac{p_{A}^{\prime}u_{A}-p_{B}^{\prime}u_{B}}{p_{A}^{\prime}+p_{B}^{\prime}}.$ (17)
The terms involving $\tilde{u}_{\mathrm{ISI}}$ represent corrections due to backscattering effects. Under normal operating conditions, these terms are small and can be calibrated out, but they are included here for completeness.
Similarly, for the ALDM interferometer, the signal at the photodiode C and D is given by
$S_{D}-S_{C}$ $=2I_{1}\left[p_{D}^{\prime 2}-p_{C}^{\prime 2}+u_{D}^{2}-u_{C}^{2}+2(u_{D}+u_{C})\sin\tilde{\psi}\right.$ (18)
$\left.+2(p_{D}^{\prime}+p_{C}^{\prime})\sin(\Omega t+\phi+\Delta\phi)+2(p_{D}^{\prime}u_{D}-p_{C}^{\prime}u_{C})\cos(\Omega t+\phi+\Delta\phi-\tilde{\psi})\right]$
where
$\Delta\phi=\Delta\phi_{\mathrm{Tx}}+\Delta\phi_{\mathrm{Rx}}+\Delta\phi_{\mathrm{PS}}$ (19)
represents the phase shift relative to the ISI. Here, $\Delta\phi_{\mathrm{Tx}}$ and $\Delta\phi_{\mathrm{Rx}}$ are the phase differences arising from the path length differences of the transmitted light and received light, respectively. $\Delta\phi_{\mathrm{PS}}$ is the phase shift introduced by the phase shifter, which can be preset on Earth. The term $I_{1}$ is the power of the reference beam in the ALDM interferometer, and $p_{\mathrm{ALDM}}^{2}=p_{C}^{2}+p_{D}^{2}$ and $u_{\mathrm{ALDM}}^{2}=u_{C}^{2}+u_{D}^{2}$ denote the ratios of the power of the received light and the backscattered light to the reference beam, respectively. Likewise, the power of the received light in the ALDM interferometer is modulated by the birefringence effect
$p_{\mathrm{ALDM}}^{\prime}=\sin\Delta\phi_{a}p_{\mathrm{ALDM}}.$ (20)
For simplicity, we assume $I_{1}=I_{0}$, which implies $p_{\mathrm{ALDM}}=p_{\mathrm{ISI}}$. In addition, the heterodyne signals are considered to be balanced Carleton and Maloney (1968); Fleddermann et al. (2018) with $p_{k}=p/\sqrt{2},\,k\in\{A,B,C,D\}$. Under these conditions, the combined AC components of the signal differences from both interferometers can be expressed as
$AC\{(S_{D}-S_{C})+(S_{B}-S_{A})\}$ (21)
$=$ $\sqrt{2}pI_{0}\Big{[}\cos\Delta\phi_{a}\sin(\Omega t+\phi)+\sin\Delta\phi_{a}\sin(\Omega t+\phi+\Delta\phi)$
$+$ $(u_{B}-u_{A})\cos(\Omega t+\phi-\psi)+(u_{D}-u_{C})\cos(\Omega t+\phi+\Delta\phi-\tilde{\psi})\Big{]}$
$=$ $\sqrt{2}pI_{0}\Big{[}\sin(\Omega t+\phi+\Delta\phi_{a})+\sin\Delta\phi_{a}[\sin(\Omega t+\phi+\Delta\phi)-\cos(\Omega t+\phi)]$
$+$ $(u_{B}-u_{A})\cos(\Omega t+\phi-\psi)+(u_{D}-u_{C})\cos(\Omega t+\phi+\Delta\phi-\tilde{\psi})\Big{]}.$
By combining the signals from both interferometers, we create a measurement that is directly sensitive to the ALDM-induced phase shift $\Delta\phi_{a}$. The algebraic manipulation above demonstrates how the combined signal contains information about both the standard interferometric phase and the ALDM-induced birefringence effect. Regulating $\Delta\phi=\pi/2$, we obtain
$AC\{(S_{D}-S_{C})+(S_{B}-S_{A})\}$ (22)
$=$ $\sqrt{2}pI_{0}\Big{[}\sin(\Omega t+\phi+\Delta\phi_{a})$
$+$ $(u_{B}-u_{A})\cos(\Omega t+\phi-\psi)-(u_{D}-u_{C})\sin(\Omega t+\phi-\tilde{\psi})\Big{]}.$
The choice of $\Delta\phi=\pi/2$ is a crucial, as it maximizes the sensitivity to the ALDM signal by ensuring that the auxiliary interferometer operates at the optimal phase difference relative to the main interferometer. This configuration converts the amplitude modulation into a phase shift that can be directly measured by standard phase detection techniques.
The phase of the heterodyne-signal can then be recovered as
$\phi_{\mathrm{ALDM}}\approx\phi+\Delta\phi_{a}+\pi/2+\phi_{b},$ (23)
where
$\phi_{b}\approx\left[(u_{A}-u_{B})\cos\psi-(u_{C}-u_{D})\sin\tilde{\psi}\right]\left[1-(u_{A}-u_{B})\sin\psi-(u_{C}-u_{D})\cos\tilde{\psi}\right].$ (24)
The phase $\phi_{\mathrm{ALDM}}$ is the key measurand of the ALDM interferometer. By subtracting $\phi_{\mathrm{ISI}}$ from $\phi_{\mathrm{ALDM}}$, we can recover the change in the polarization angle induced by the ALDM for a single-link light, as expressed in Eq. (8).
In the subsequent analysis, we assume that the contribution of backscattered light is negligible. This assumption is reasonable, as modern space-based interferometer designs incorporate advanced stray light suppression techniques. Moreover, the influence of backscattered light can be further mitigated through meticulous optical design and sophisticated signal processing methods.
IV Sensitivity to the ALDM-photon coupling
In this section, we derive the sensitivity of the modified SGWDs to the ALDM-photon coupling $g_{a\gamma}$. We begin by analyzing the single-arm signal and subsequently extend the analysis to various TDI combinations.
IV.1 Single-arm signal
By differentiating the phase difference induced by the ALDM-induced birefringence, we define the single-arm response as  [^1] The single-arm signal seems a little different from the equation in Gué et al. (2025), this originates from the typo of the definition of the reduced Planck energy in their paper, which is supposed to be $E_{P}=\sqrt{\hbar c/\kappa}\simeq 2.435\times 10^{18}\,\mathrm{GeV}$. With the correct definition, the factor $\sqrt{8\pi G}$ would be canceled out.
$\eta_{rs}$ $\equiv\frac{1}{\omega_{r}}\frac{d\left(\phi_{\mathrm{ALDM}}-\phi_{\mathrm{ISI}}\right)}{dt}\approx\frac{1}{\omega_{r}}\frac{d\Delta\phi_{a}}{dt}$ (25)
$=-\frac{g_{a\gamma}}{\omega_{r}}\sqrt{2\rho_{\mathrm{DM}}}\sin\frac{m_{a}L}{2}\cos\Big{(}m_{a}t+\Phi-\frac{m_{a}L}{2}\Big{)},$
and similarly
$\bar{\eta}_{rs}\equiv\frac{1}{\omega_{r}}\frac{d\left(\phi_{\mathrm{ISI}}-\phi_{\mathrm{ALDM}}\right)}{dt}\approx-\frac{1}{\omega_{r}}\frac{d\Delta\phi_{a}}{dt},$ (26)
where $\omega_{r}$ denotes the laser frequency, and the subscripts $r$ and $s$ represent the indices of the S/C that received and sent the light beam, respectively. The single-arm responses $\eta_{rs}$ and $\bar{\eta}_{rs}$ provide a direct measure of the time derivative of the ALDM-induced phase shift. In practice, SGWDs employ TDI to suppress noises, such as laser frequency noise and OB motion noise, by combining the signals from multiple interferometric arms with appropriate time delays Armstrong et al. (1999); Tinto and Armstrong (1999); Dhurandhar et al. (2002); Tinto and Dhurandhar (2021). This technique enhances sensitivity to gravitational wave signals. From the definition of the single-arm response to the ALDM effect, the subtraction between the two channels cancels out these common-mode noises.
Working with the time derivative of the phase difference, rather than the phase difference itself, offers significant advantages for signal processing. It facilitates the implementation of TDI techniques and enable the separation of the ALDM signal from slowly varying systematic effects. Additionally, the frequency dependence of the signal provides a distinctive signature, aiding in the discrimination against potential background sources.
IV.2 Signals for different TDI combinations
TDI is implemented in SGWDs to synthesize virtual interferometric channels immune to dominant noise sources while preserving the physical signals of interest. This technique is essential for achieving the sensitivity required for gravitational wave detection, and we can leverage the same approach for ALDM searches without compromising the primary science objectives.
For each TDI combination, the ALDM-induced signal is weighted by a factor that depends on the arm length $L$ and the ALDM mass $m_{a}$. For example, the Sagnac variable $\alpha$ is given by Armstrong et al. (1999)
$\alpha(t)$ $=\Big{[}\bar{\eta}_{12}+D_{12}\bar{\eta}_{23}+D_{12}D_{23}\bar{\eta}_{31}\Big{]}-\Big{[}\eta_{13}+D_{13}\eta_{32}+D_{13}D_{32}\eta_{21}\Big{]}$ (27)
$=-2\Big{(}\eta_{13}+D_{13}\eta_{32}+D_{13}D_{32}\eta_{21}\Big{)}$
$=2\,\frac{g_{a\gamma}}{\omega_{r}}\sqrt{2\rho_{\mathrm{DM}}}\sin\frac{3m_{a}L}{2}\cos\Big{(}m_{a}t+\Phi-\frac{3m_{a}L}{2}\Big{)},$
where $D_{ij}$ represents the delay operator corresponding to the light travel time between S/C $i$ and $j$. Extending this to the second-generation Sagnac variable Vallisneri (2005), we obtain:
$\alpha_{2}(t)$ $=\Big{(}1-D_{12}D_{23}D_{31}\Big{)}\alpha$ (28)
$=4\,\frac{g_{a\gamma}}{\omega_{r}}\sqrt{2\rho_{\mathrm{DM}}}\sin^{2}\frac{3m_{a}L}{2}\sin\Big{(}m_{a}t+\Phi-\frac{3m_{a}L}{2}\Big{)}.$
Unlike the modifications proposed in Gué et al. (2025); Yao et al. (2025), our approach not only leverages the Sagnac $\alpha$ combination but is also applicable to other TDI combinations. For example, the X variable reads
$X(t)$ $=\Big{(}1-D_{13}D_{31}\Big{)}\Big{(}\eta_{12}+D_{12}\eta_{21}\Big{)}-\Big{(}1-D_{12}D_{21}\Big{)}\Big{(}\bar{\eta}_{13}+D_{13}\bar{\eta}_{31}\Big{)}$ (29)
$=4\,\frac{g_{a\gamma}}{\omega_{r}}\sqrt{2\rho_{\mathrm{DM}}}\sin^{2}(m_{a}L)\sin\Big{(}m_{a}t+\Phi-2m_{a}L\Big{)},$
with its second-generation combination
$X_{2}(t)$ $=\Big{(}1-D_{31}^{2}D_{12}^{2}\Big{)}X$ (30)
$=8\,\frac{g_{a\gamma}}{\omega_{r}}\sqrt{2\rho_{\mathrm{DM}}}\sin^{2}(m_{a}L)\sin(2m_{a}L)\cos\Big{(}m_{a}t+\Phi-4m_{a}L\Big{)}.$
The fully symmetric $\zeta$ variable is expressed as Muratore et al. (2022)
$\zeta(t)$ $=D_{12}\Big{(}\bar{\eta}_{31}-\eta_{32}\Big{)}+D_{23}\Big{(}\bar{\eta}_{12}-\eta_{13}\Big{)}+D_{31}\Big{(}\bar{\eta}_{23}-\eta_{21}\Big{)}$ (31)
$=6\,\frac{g_{a\gamma}}{\omega_{r}}\sqrt{2\rho_{\mathrm{DM}}}\sin\frac{m_{a}L}{2}\cos\Big{(}m_{a}t+\Phi-\frac{3m_{a}L}{2}\Big{)},$
with a related second-generation variable Muratore et al. (2020)
$\zeta_{2}(t)$ $=\Big{(}D_{31}-D_{12}D_{23}\Big{)}\zeta$ (32)
$=-12\,\frac{g_{a\gamma}}{\omega_{r}}\sqrt{2\rho_{\mathrm{DM}}}\sin^{2}\frac{m_{a}L}{2}\cos\Big{(}m_{a}t+\Phi-3m_{a}L\Big{)}.$
For a signal $a(t)=A\sin(m_{a}t+\phi)$ or $a(t)=A\cos(m_{a}t+\phi)$, its one-sided power spectral density (PSD) is given by
$S(f)=\frac{A^{2}}{2\pi^{2}T_{obs}}\frac{\sin^{2}[\pi(f-f_{a})T_{obs}]}{(f-f_{a})^{2}},$ (33)
where $f_{a}=m_{a}/2\pi$. For each TDI variable $i\in\{\alpha,\alpha_{2},X,X_{2},\zeta,\zeta_{2}\}$, the one-sided PSD of the signal is
$S_{i}(f)=\frac{4C_{i}^{2}(f)g_{a\gamma}^{2}\rho_{\mathrm{DM}}}{\pi^{2}\omega_{r}^{2}T_{\mathrm{obs}}}\frac{\sin^{2}[\pi(f-f_{a})T_{\mathrm{obs}}]}{(f-f_{a})^{2}},$ (34)
where $f_{a}=m_{a}/2\pi$ is the ALDM Compton frequency, and $T_{\mathrm{obs}}$ is the observation time. The factor $C_{i}(f)$ is a sine term that encapsulates the response of the specific TDI combination
$C_{X}(f)=2\sin^{2}(2\pi fL),\quad C_{X_{2}}(f)=4\sin^{2}(2\pi fL)\sin(4\pi fL),$ (35)
$C_{\alpha}(f)=\sin(3\pi fL),\quad C_{\alpha_{2}}(f)=2\sin^{2}(3\pi fL),$ (36)
$C_{\zeta}(f)=3\sin(\pi fL),\quad C_{\zeta_{2}}(f)=6\sin^{2}(\pi fL).$ (37)
In the limit of $f\to f_{a}$, Eq. (34) simplifies to
$S_{i}(f)=4C_{i}^{2}(f)g_{a\gamma}^{2}\rho_{\mathrm{DM}}T_{\mathrm{obs}}/\omega_{r}^{2}.$ (38)
IV.3 Expected sensitivity
The two primary noise sources that limit the sensitivity of SGWDs are optical metrology system (OMS) noise and test mass acceleration noise. The OMS noise arises from imperfections in the laser interferometry system used to measure the distance between test masses in different spacecraft, such as residual frequency noise, shot noise, and pathlength noise. The acceleration noise originates from non-gravitational forces acting on the test masses. In our analysis, we assume that the proposed modifications to the OB do not significantly alter the noise PSDs. For completeness, we note that the noise PSDs for the $X$, $\alpha$, and $\zeta$ channels, as well as their second-generation counterparts, are given by expressions of the form Hartwig et al. (2023) [^2] Here we assume that the OMS and acceleration contributions for the Doppler variables $\eta$ and $\bar{\eta}$ are uncorrelated.
$N_{X}(f)$ $=16\sin^{2}(2\pi fL)\{[3+\cos(4\pi fL)]S_{\mathrm{acc}}(f)+S_{\mathrm{oms}}(f)\},$ (39)
$N_{X_{2}}(f)$ $=4\sin^{2}(4\pi fL)\,N_{X}(f),$ (40)
$N_{\alpha}(f)$ $=8\Big{[}2\sin^{2}(\pi fL)+\sin^{2}(3\pi fL)\Big{]}S_{\mathrm{acc}}(f)+6S_{\mathrm{oms}}(f),$ (41)
$N_{\alpha_{2}}(f)$ $=4\sin^{2}(3\pi fL)\,N_{\alpha}(f),$ (42)
$N_{\zeta}(f)$ $=6\Big{[}4\sin^{2}(\pi fL)S_{\mathrm{acc}}(f)+S_{\mathrm{oms}}(f)\Big{]},$ (43)
$N_{\zeta_{2}}(f)$ $=4\sin^{2}(\pi fL)\,N_{\zeta}(f).$ (44)
Here, $S_{\mathrm{oms}}(f)$ and $S_{\mathrm{acc}}(f)$ denote the PSDs of the OMS and acceleration noises, respectively. For LISA, TianQin and Taji, these noise PSDs are given by Babak et al. (2021)
$S_{\mathrm{oms}}(f)$ $=\left(2\pi fP_{\mathrm{oms}}/c\right)^{2}\left[1+\left(\frac{2\times 10^{-3}\>\mathrm{Hz}}{f}\right)^{4}\right]\>\mathrm{Hz}^{-1}\,,$ (45)
$S_{\mathrm{acc}}(f)$ $=\left(\frac{P_{\mathrm{acc}}}{2\pi fc}\right)^{2}\left[1+\left(\frac{0.4\times 10^{-3}\>\mathrm{Hz}}{f}\right)^{2}\right]\,\left[1+\left(\frac{f}{8\times 10^{-3}\>\mathrm{Hz}}\right)^{4}\right]\>\mathrm{Hz}^{-1}\,,$
where $P_{\mathrm{oms}}$ and $P_{\mathrm{acc}}$ are the noise amplitude parameters. For the BBO, the noises PSDs are given by Corbin and Cornish (2006)
$S_{\mathrm{oms}}^{\mathrm{BBO}}(f)$ $=\frac{2.0\times 10^{-34}}{3L^{2}}\>\mathrm{m^{2}\cdot Hz^{-1}},$ (46)
$S_{\mathrm{acc}}^{\mathrm{BBO}}(f)$ $=\frac{4.5\times 10^{-34}}{(2\pi f)^{4}(3L)^{2}}\>\mathrm{m^{2}\cdot Hz^{4}\cdot Hz^{-1}}.$
Parameter LISA Taiji TianQin BBO
Arm $\text{Arm length}~{}L~{}(10^{9}\mathrm{m})$ 2.5 3 0.17 0.05
$P_{\mathrm{oms}}~{}(10^{-12}\mathrm{m})$ 15 8 1 /
$P_{\mathrm{acc}}~{}(10^{-15}\mathrm{m}\cdot\mathrm{s}^{-2})$ 3 3 1 /
Laser frequency $\nu_{0}~{}(\mathrm{Hz})$ $2.82\times 10^{14}$ $2.82\times 10^{14}$ $2.82\times 10^{14}$ $6\times 10^{14}$
Integration time $T_{\mathrm{obs}}~{}(\mathrm{year})$ $4.5$ $5$ $5/2$ $4$
Frequency band $(\mathrm{Hz})$ $[10^{-4},1]$ $[10^{-4},1]$ $[10^{-4},1]$ $[10^{-1},10]$
Table 1: Parameters of of several planned SGWDs Colpi et al. (2024); Luo et al. (2020, 2016); Crowder and Cornish (2005).
Assuming that all optical benches operate with the same laser frequency $\omega_{r}=2\pi\nu_{0}$, the sensitivity to the ALDM coupling can be estimated from the signal PSD as
$g_{a\gamma}^{i}(f)=\sqrt{\frac{S_{i}(f)\,\pi^{2}\nu_{0}^{2}}{\rho_{\mathrm{DM}}\,C_{i}^{2}(f)\,T_{\mathrm{obs}}}}$ (47)
The signal-to-noise ratio (SNR) for each channel is defined by
$\mathrm{SNR}=\frac{S_{i}}{N_{i}}.$ (48)
We assume an SNR of 1 to provide a conservative estimate of the interferometer’s performance. It is noteworthy that the ratio of noise PSD between the first- and second-generation TDI combinations is the same as the ratio of the squares of their corresponding factor $C_{i}(f)$. Therefore, the second-generation TDI combinations ($X_{2},\,\alpha_{2},\zeta_{2}$) yield sensitivity estimates that are identical to those of their first-generation counterparts. In Fig. 2, we illustrate the sensitivity estimates of LISA, Taiji, TianQin, and BBO to the ALDM-photon coupling $g_{a\gamma}$ for the first-generation TDI combinations. These estimates are compared with existing constraints derived from CAST Anastassopoulos et al. (2017), SN 1987A Payez et al. (2015), the M87 galaxy Marsh et al. (2017) and the quasar H1821+643 Reynés et al. (2021). The parameters of interest for the sensitivity curves of various SGWDs are summarized in Table. 1.
(a) X channel
(b) $\alpha$ channel
(c) $\zeta$ channel
Figure 2: Sensitivity estimates of various SGWDs to the ALDM-photon coupling $g_{a\gamma}$ for the first-generation TDI conbinations. The constraints from CAST Anastassopoulos et al. (2017), SN 1987A Payez et al. (2015), the M87 galaxy Marsh et al. (2017) and the quasar H1821+643 Reynés et al. (2021) are also shown for comparison.
The comparisons to existing astrophysical and laboratory constraints reveal that the SGWD-based approach holds potential to improve sensitivities to the ALDM-photon coupling strength, particularly in the mass range of $10^{-19}\,\mathrm{eV}$ to $10^{-14}\,\mathrm{eV}$. Notably, the SGWD senstitivity curves dip below the astrophysical constraints in certain mass windows, reaching $g_{a\gamma}\sim 10^{-13}\>\mathrm{GeV^{-1}}$, signify their competitive reach in the search for ALDM.
The sensitivity curves for various experimental configurations and TDI channels demonstrate that these detectors can probe a wide range of ALDM masses. The inclusion of different SGWDs shows their complementary coverage across the ALDM parameter space. The BBO extends sensitivity toward higher-frequency domains, covering a distinct region of parameter space. The sensitivity curves also indicate that certain TDI combinations offer superior detection potential for specific mass ranges. In particular, the $\zeta$ channel exhibits better performance across a wide range of ALDM masses.
Future work could refine these projections by incorporating potential noise sources and optimizing interferometric configurations to further enhance detection capabilities.
V Discussion
The proposed OB modification offers a promising route to probe the ALDM-photon coupling without altering the polarization state of inter-spacecraft links. However, the practical implementation of this modification entails a series of intricate engineering challenges that demand meticulous consideration. For instance, the stabilization of additional optical elements is critical to avoid introducing extra phase noise. Specifically, the 50/50 beam splitter ratio for the ALDM interferometer must be carefully maintained to prevent noise arising from power fluctuations, while additional baffling for the auxiliary optical path is required to mitigate stray light effects.
The projected sensitivity curves, presented in the preceding section, provide preliminary assessment of the instrument’s performance, demonstrating the potential of SGWDs to probe ALDM. Despite these promising sensitivity projections, the impact of the modification on the overall OB design cannot be overlooked. The integration of an auxiliary interferometer into the OB introduces challenges not only in maintaining mechanical and thermal stability but also in managing the total light power budget. Systematic errors arising from the auxiliary path, such as those induced by imperfect beam splitting or misalignment, must be carefully quantified and minimized. The p-polarized light induced by the ALDM may be mitigated by the the existing optics of the OB like transimitting beam clipping, requiring dedicated design for the actual light path. Future research should focus on detailed simulations and laboratory demonstrations to validate the preliminary sensitivity estimates and to develop practical methods for mitigating the additional noise contributions introduced by the modification.
In summary, while the analytical projections and sensitivity curves are encouraging, the operational realization of the proposed OB modification in SGWDs necessitates the resolution of significant engineering hurdles. The feasibility of this approach hinges on the successful integration of the auxiliary interferometer without compromising the performance of the primary gravitational wave detection channels. If these challenges can be effectively addressed, the modified OB design could emerge as a powerful new tool in the search for ALDM, thereby expanding the scientific capabilities of SGWDs.
VI Conclusion
In this work, we have proposed a modification to the OB in SGWDs to detect ALDM. We presented an alternative strategy by introducing an auxiliary interferometer that converts ALDM-induced polarization rotations into measurable phase shifts, without changing the polarization of the inter-spacecraft laser light. The projected sensitivity to the ALDM-photon coupling covers a broad range of ALDM masses, from $10^{-19}\,\mathrm{eV}$ to $10^{-14}\,\mathrm{eV}$. In general, SGWDs such as LISA, Taiji, and TianQin could achieve sensitivities of $g_{a\gamma}\sim 10^{-13}\>\mathrm{GeV^{-1}}$ for ALDM with masses around $10^{-17}\>\mathrm{eV}$. A combined analysis of all TDI channels could further improve the reach. This approach preserves the design integrity of current SGWD optical benches while significantly expanding their scientific capability to explore new physics.
Acknowledgements.
QGH is supported by the National Key Research and Development Program of China Grant No.2020YFC2201502, grants from NSFC (grant No. 12475065, 12250010). RMY is supported by the National Natural Science Foundation of China under Grant No. 12447131. BXJ is supported by the National Natural Science Foundation of China under Grant No. 12447105.